# ROLCH: Regularized Online Learning for Conditional Heteroskedasticity

### v1

@@ 2.2 Online estimation of regularized linear models
      Check convergence for $\widehat{\boldsymbol{\beta}}_{n+1,\lambda}$ and proceed to next $\lambda$ if converged.
Return: $\widehat{\boldsymbol{\beta}}_{n+1}$ for all $\lambda$
Algorithm 1 Online LASSO
@@
To improve the convergence speed in estimating the full $\lambda$-path, the $\boldsymbol{\beta}_{\lambda}$ for the previous regularization parameter is used as the start value. Also, we can employ potentially weighted Gramian matrices ${\mathcal{G}}={\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{X}}_{n}$ and ${\mathcal{H}}={\mathcal{X}}^{\top}_{n}{\mathcal{W}}_{n}{\mathcal{Y}}_{n}$ and the respectively exponentially discounted versions of ${\mathcal{G}}_{n}$ and ${\mathcal{H}}_{n}$ as in the previous paragraphs (Messner & Pinson, 2019). Our implementation in the software package ROLCH exploits active set iterations to speed up convergence, i.e. after the first full cycle through all features $j$, we only update non-zero coefficients (Meier et al., 2008; Friedman et al., 2010). A further overview on different implementations, tips and tricks can be found in Shi et al. (2016).
Online LARS
The original LARS algorithm introduced by Efron et al. (2004) is already using the Gramian matrix, so the adaption to an online setting is straightforward. Maheshwari et al. (2017) have also proposed a LARS-like online learning algorithm. However, compared to the OCD algorithm for the LASSO estimator, we cannot warm-start the online LARS implementation. Therefore, we prefer the LASSO estimator throughout the remainder of the paper.
@@
We select the regularization within the iterative re-weighting loop using information criteria. We define the generalised information criterion (GIC) as
$\operatorname{GIC}(\hat{{\mathcal{L}}},\nu)=-2\log(\hat{{\mathcal{L}}})+\nu_{0}k+\nu_{1}k\log(n)+\nu_{2}k\log(\log(n))$ (24)
@@
where $k$ denotes the number of estimated parameters, $\nu=(\nu_{0},\nu_{1},\nu_{2})$ denotes a triplet of parameters, and $\hat{{\mathcal{L}}}$ denotes the maximized value of the likelihood function of the model. The most commonly used information criteria Akaike’s Information Criterion (AIC, $\nu=(2,0,0)$), Bayesian Information Criterion (BIC, $\nu=(0,1,0)$) and the Hamann-Quinn Criterion (HQC, $\nu=(0,0,2)$) can be derived of the GIC and are implemented in our software package as tuning hyper-parameters Ding et al. (2018). Under the Gaussian assumption, the likelihood can be formulated as a function of the residual sum-of-squares
$\log(\hat{{\mathcal{L}}})=-\frac{n}{2}+\log\left(\frac{\text{RSS}}{n}\right)+C$ (25)
@@
where $C=-\frac{n}{2}\left(1+\log(2\pi)\right)$ is a constant which only depends on the data and hence can be neglected if the data underlying the model selection is the same for all models. This allows an efficient online update of the information criterion, including the option to account for exponential forget. Note that we use the effective training length $n=\frac{(1-\gamma^{N})}{(1-\gamma)},$ where $N$ is the number of data points seen by the model (so far). Alternatively, Messner & Pinson (2019) propose to use the exponentially discounted in-sample root mean square error to select the optimal $\lambda$ essentially taking $\frac{1}{n}\text{RSS}$. However, this approach will converge towards the OLS solution.
2.3 Online GAMLSS
@@
Having established the batch RS algorithm and the online coordinate descent for regularized linear models, this section proceeds by putting the pieces together and presenting the online linear parametric GAMLSS model.
Method
@@
We aim to present the algorithm in the most generic form here. After having seen $n$ observations, we have estimated coefficients $\widehat{\boldsymbol{\beta}}_{k,n}$ for $k=1,..,p$ distribution parameters. Given some new data $\boldsymbol{Y}_{n+1}$ for the response and $\boldsymbol{X}_{k,n+1}$ covariate design matrices, we aim to update our coefficients towards $\widehat{\boldsymbol{\beta}}_{k,n+1}$ for all $k=1,...,p$. Let us recall the definition and update rule for the weighted, exponentially discounted Gramian matrices ${\mathcal{G}}_{k,n}={\mathcal{X}}^{\top}_{k,n}\Gamma_{k,n}{\mathcal{W}}_{k,n}{\mathcal{X}}_{k,n}$ and ${\mathcal{H}}_{k,n}={\mathcal{X}}^{\top}_{k,n}\Gamma_{k,n}{\mathcal{W}}_{k,n}{\mathcal{Y}}_{k,n}$ here, to which we will refer throughout the algorithm.
${\mathcal{G}}_{k,n+1}$ $=\gamma{\mathcal{G}}_{k,n}+w_{k,n+1}(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{X}_{k,n+1})$ (26)
${\mathcal{H}}_{k,n+1}$ $=\gamma{\mathcal{H}}_{k,n}+w_{k,n+1}(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{Y}_{k,n+1})$ (27)
@@
On a high level, for updating the coefficients in the online GAMLSS, we do the following steps in the outer cycle: we initialise the $\boldsymbol{\eta}_{k,n+1}^{[0,0]}=g_{k}\left(\boldsymbol{\beta}_{k,n}\boldsymbol{X}^{\top}_{k,n+1}\right)$ and this allows us to evaluate $\boldsymbol{u}_{k,n+1}^{[0,0]}$, $\boldsymbol{W}_{kk,n+1}^{[0,0]}$ and $\boldsymbol{z}_{k,n+1}^{[0,0]}$. From there, we can update the weighted, exponentially discounted Gramian matrices ${\mathcal{G}}_{k,n}\rightarrow{\mathcal{G}}_{k,n+1}^{[0,0]}$ and ${\mathcal{H}}_{k,n}\rightarrow{\mathcal{H}}_{k,n+1}^{[0,0]}$ and subsequently run the online coordinate descent (OCD) algorithm to update the coefficient path in the inner iteration. As in the batch algorithm, we run the outer and inner cycles until convergence.
Detailed Update Algorithm
The following paragraph describes the update step for the algorithm in detail. We keep the structure from Section 2.1. To make the discussion clearer, we call an update step a full update for all distribution parameters $\theta_{k,n}\rightarrow\theta_{k,n+1}$ and associated regression coefficients $\boldsymbol{\beta}_{k,n}\rightarrow\boldsymbol{\beta}_{k,n+1}$ for $k=1,...,p$ while after receiving data $\boldsymbol{Y}_{k,n+1}$ and $\boldsymbol{X}_{k,n+1}$, while step might refer to any arbitrary step in the algorithm. The update step does the following:
  1. 1.
    (Initialisation:) Initialise the fitted values $\theta_{k,n+1}^{[0,0]}=\boldsymbol{\beta}_{k,n}\boldsymbol{X}^{\top}_{k,n+1}$ for $k=1,...,p$. Evaluate the linear predictors ${\eta^{[0,0]}_{k,n+1}=g_{k}(\theta_{k,n+1}^{[0,0]})}$ for $k=0,...,p$. Note that for a one-step update step, these values are scalars.
  2. 2.
    (Outer Cycle:) Start the outer cycle $m=0,1,...$ until convergence. Iterate over $k=1,....,p$ and
    1. (a)
      Start the inner cycle $n=0,1,...$ until convergence:
      • i.
        Evaluate $u_{n+1,k}^{[m,r]}$, $w_{n+1,kk}^{[m,r]}$ and $z_{n+1,k}^{[m,r]}$ using Equations 4, 5 and 6.
      • ii.
@@
        Update the Gramian Matrices ${\mathcal{G}}_{k,n}\rightarrow{\mathcal{G}}_{k,n+1}^{[m,r]}$ and ${\mathcal{H}}_{k,n}\rightarrow{\mathcal{H}}_{k,n+1}^{[m,r]}$ by taking the working vector $z_{n+1,k}^{[m,r]}$ as response variable and weights $w_{kk,n+1}^{[m,r]}$:
        ${\mathcal{G}}_{k,n+1}^{[m,r]}$ $=\gamma{\mathcal{G}}_{k,n}+w_{kk,n+1}^{[m,r]}\left(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{X}_{k,n+1}\right)$
        ${\mathcal{H}}_{k,n+1}^{[m,r]}$ $=\gamma{\mathcal{H}}_{k,n}+w_{kk,n+1}^{[m,r]}\left(\boldsymbol{X}^{\top}_{k,n+1}z_{n+1,k}^{[m,r]}\right)$
      • iii.
@@
        Update the coefficient path $\boldsymbol{\beta}_{n,k,\lambda}^{[m,r+1]}\rightarrow\boldsymbol{\beta}_{n,k,\lambda}^{[m,r+1]}$ using OCD based on ${\mathcal{G}}_{k,n+1}^{[m,r]}$ and ${\mathcal{H}}_{k,n+1}^{[m,r]}$ and select the optimal tuning parameter $\lambda$ using information criteria.
        • A.
          Update the exponentially discounted weighted sum of squared residuals to approximate the likelihood (see Equation 25) by taking
          $\text{RSS}_{k,n+1,\lambda}=\dfrac{\left(w_{kk,n+1}\left(z_{n+1,k}^{[m,r]}-\boldsymbol{\beta}_{n,k,\lambda}^{[m,r+1]}\boldsymbol{X}^{\top}_{k,n+1}\right)\right)^{2}+(1-\gamma)\text{RSS}_{n,k,\lambda}}{w_{kk,n+1}+(1-\gamma)\omega_{kk,n}}$ (28)
          where $\omega_{kk,n}$ is the sum of discounted weights $\omega_{kk,n}=\sum_{i=0}^{n}\gamma^{i-1}w_{kk,i}$ up to $n$, which can easily be tracked online.
        • B.
          Select the $\lambda^{\text{opt}}_{k,n+1}=\operatorname{arg\;min}_{\lambda}\left(\operatorname{GIC}(\hat{{\mathcal{L}}}_{\lambda},\nu)\right)$ and update the likelihood based on the RSS by Equation 25. Select the tuned coefficients $\boldsymbol{\beta}_{n+1,k}^{[m,r+1]}=\boldsymbol{\beta}_{n+1,k,\lambda^{\text{opt}}}^{[m,r+1]}$.
      • iv.
        Calculate the updated $\boldsymbol{\eta}_{k,n+1}^{[m,r+1]}$ and $\boldsymbol{\theta}_{k,n+1}^{[m,r+1]}$.
      • v.
        Evaluate the convergence.
    2. (b)
      End the inner cycle on the convergence of $\boldsymbol{\beta}_{k,n+1}^{[m,...]}$ and set $\boldsymbol{\beta}_{k,n+1}^{[m+1,0]}=\boldsymbol{\beta}_{k,n+1}^{[m,...]}$, set $\boldsymbol{\eta}_{k,n+1}^{[m+1,0]}=\boldsymbol{\eta}_{k,n+1}^{[m,...]}$ and set $\boldsymbol{\theta}_{k,n+1}^{[m+1,0]}=\boldsymbol{\theta}_{k,n+1}^{[m,...]}$. Otherwise, continue the inner cycle until convergence.
  3. 3.
    (Next Parameter:) Update $k$ upon convergence.
  4. 4.
    (End:) End the outer cycle if the change in the penalized likelihood is sufficiently small, otherwise update $m$ and continue the outer cycle.
@@
Note that within step 2 – a, in each inner iteration, we start at the Gramian matrices of the previous full fit, not at the previous iterations $m$ or $r$ of the algorithm, since this would imply adding the $\boldsymbol{X}_{k,n+1}$ and $z_{n+1,k}^{[m,r]}$ multiple times to the Gramian matrices within one update step. However, we can (and should) warm-start the OCD algorithm using the coefficient path from the previous iterations within each inner iteration of the update step.
@@
The main convergence criteria is the difference between the log-likelihood under the distribution $\mathcal{D}$, given the inner and outer iterations current fitted values $\hat{\boldsymbol{\eta}}^{[m,r]}_{k,n+1}$, both for the batch and online case. We track the (exponentially discounted) log-likelihood for checking the convergence in the online case. Additionally, we introduce a breaking criterion if the residual sum of squares increases during the inner cycle by more than a factor of $\varepsilon_{\text{RSS}}=1.5$ of the previous inner cycle’s RSS. This criterion can be turned off by setting $\varepsilon_{\text{RSS}}$ to an arbitrarily high number or infinity. This criterion can be used to ’force’ the algorithm into a local minima with respect to the RSS in the inner cycle, compared to global minima in the deviance. This can improve the convergence, if the parametric distribution assumption does not fit the observed data well, especially in the case of outliers.
Mini-batch update steps
The algorithm above implements one-step updates, i.e. receiving new data row-by-row: in each update, we receive $\boldsymbol{X}_{k,n+1}$ and $\boldsymbol{Y}_{k,n+1}$. In principle, mini-batch update steps, in which we receive ${\mathcal{X}}_{k,n+1:n+t}=(\boldsymbol{X}_{k,n+1},...,\boldsymbol{X}_{k,n+t})^{\top}$ and ${\mathcal{Y}}_{n+1:n+t}=(Y_{n+1},...,Y_{n+t})$ in a single pass can be implemented in the same manner. Equations 26 and 27 natively handle mini-batch updates. Equation 28 needs to be adjusted to
$\text{RSS}_{k,n+t,\lambda}=\dfrac{\left((1-\gamma)^{t-1}\boldsymbol{W}_{kk,n+t}\left(\boldsymbol{z}_{n+t,k}^{[m,r]}-\boldsymbol{\beta}_{n+t,k,\lambda}^{[m,r]}{\mathcal{X}}^{\top}_{k,n+1:n+t}\right)\right)^{2}+(1-\gamma)^{t}\text{RSS}_{n,k,\lambda}}{\sum_{i=0}^{t-1}(1-\gamma)^{i}w_{kk,n+i}+(1-\gamma)^{t}\omega_{kk,n}}$
@@
    For the selection of the regularization, we suggest selecting a higher regularization for higher $k$ to avoid overfitting in modelling the conditional scale, kurtosis and skewness (Ziel, 2022). We also note that the $\varepsilon_{\lambda,k}$ for higher moments needs to be smaller, otherwise the lower end of the coefficient path might not approach the OLS solution. Groll et al. (2019) analyse the impact of different shrinkage parameters on batch distributional regression.
Limitations
@@
Finally, let us note an important limitation of our proposed algorithm. In the batch setting with repeated fits of the GAMLSS, the weight matrix ${\mathcal{W}}_{kk}^{[m,r]}=\operatorname{diag}\left(w_{kk,0},...,w_{kk,n}\right)$ is updated in every iteration $m,r$ and again, in the next batch fit for $n+1$, the weight matrix is updated for all for $w_{kk,i}$ from $i=0,...,n+1$ in all iterations. In the online setting, ${\mathcal{W}}_{kk,n}$ is the weight matrix after convergence of update step $n-1\rightarrow n$. In the update step $n\rightarrow n+1$, we cannot update ${\mathcal{W}}_{kk,n}$ anymore. Therefore, we set
${\mathcal{W}}_{kk,n+1}^{[m,r]}=\begin{pmatrix}(1-\gamma){\mathcal{W}}_{kk,n}&0\\ 0&w_{kk,n+1}^{[m,r]}\end{pmatrix}$

### v2

@@ 7                   Update $\hat{\beta}_{j,\lambda}$ according to Equation 19
9            Check convergence for $\widehat{\boldsymbol{\beta}}_{n+1,\lambda}$ and proceed to next $\lambda$ if converged.
10
11
Output: $\widehat{\boldsymbol{\beta}}_{n+1}=\left(\hat{\beta}_{j,\lambda},...\right)^{\top}$ for all $\lambda\in\boldsymbol{\lambda}$
Algorithm 3 Online LASSO, see Angelosante et al. (2010) and Messner and Pinson (2019)
@@
To improve the convergence speed in estimating the full $\lambda$-path, the $\boldsymbol{\beta}_{\lambda}$ for the previous regularization parameter is used as the start value. Our implementation in the software package ROLCH exploits active set iterations to speed up convergence, i.e. after the first full cycle through all features $j$, we only update non-zero coefficients (Meier et al., 2008; Friedman et al., 2010). A further overview on different implementations, tips and tricks can be found in Shi et al. (2016).
@@
The proposed estimation on a grid of regularization parameters leads to the issue of online model selection. We propose the use of information criteria (IC) for this task. We define the generalised information criterion (GIC) as
$\operatorname{GIC}(\hat{{\mathcal{L}}},\nu)=-2\log(\hat{{\mathcal{L}}})+\nu_{0}k+\nu_{1}k\log(\tilde{N})+\nu_{2}k\log(\log(\tilde{N}))$ (20)
@@
where $k$ denotes the number of estimated parameters, $\nu=(\nu_{0},\nu_{1},\nu_{2})$ denotes a triplet of parameters, and $\hat{{\mathcal{L}}}$ denotes the maximized value of the likelihood function of the model. The most commonly used information criteria Akaike’s Information Criterion (AIC, $\nu=(2,0,0)$), Bayesian Information Criterion (BIC, $\nu=(0,1,0)$) and the Hamann-Quinn Criterion (HQC, $\nu=(0,0,2)$) can be recovered from the GIC and Kim et al. (2012); Kock (2016). Under the Gaussian assumption, the likelihood can be formulated as a function of the residual sum-of-squares
$\log(\hat{{\mathcal{L}}})=-\frac{\tilde{N}}{2}\log\left(\frac{\text{RSS}}{\tilde{N}}\right)+C$ (21)
@@
where $C=-\frac{\tilde{N}}{2}\left(1+\log(2\pi)\right)$ is a constant which only depends on the data and hence can be neglected if the data underlying the model selection is the same for all models. This allows an efficient online update of the information criterion, including the option to account for exponential forget. Note that we use the effective training length $\tilde{N}=\frac{(1-\gamma^{N})}{1-\gamma},$ where $N$ is the number of observations received so far. Alternatively, Messner and Pinson (2019) propose to use the exponentially discounted in-sample root mean square error to select the optimal $\lambda$ essentially taking $\frac{1}{\tilde{N}}\text{RSS}$. However, this approach will converge towards the OLS solution.
3 Online Distributional Regression
@@
Having established the batch GAMLSS algorithm and the online coordinate descent for regularized linear models, this section proceeds by putting the pieces together and presents the online linear parametric GAMLSS model in Algorithm 4.
Overview
@@
After having seen $n$ observations, we have estimated coefficients $\widehat{\boldsymbol{\beta}}_{k,n}$ for $k=1,..,p$ distribution parameters. Given some new data $\boldsymbol{Y}_{n+1}$ for the response and $\boldsymbol{X}_{k,n+1}$ covariate design matrices, we aim to update our coefficients towards $\widehat{\boldsymbol{\beta}}_{k,n+1}$ for all $k=1,...,p$. Let us recall the definition and update rule for the weighted, exponentially discounted Gramian matrices ${\mathcal{G}}_{k,n}={\mathcal{X}}^{\top}_{k,n}\Gamma_{k,n}{\mathcal{W}}_{k,k,n}{\mathcal{X}}_{k,n}$ and ${\mathcal{H}}_{k,n}={\mathcal{X}}^{\top}_{k,n}\Gamma_{k,n}{\mathcal{W}}_{k,k,n}{\mathcal{Y}}_{k,n}$ here, to which we will refer throughout the algorithm.
${\mathcal{G}}_{k,n+1}$ $=\gamma{\mathcal{G}}_{k,n}+w_{k,k,n+1}(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{X}_{k,n+1})$ (22)
${\mathcal{H}}_{k,n+1}$ $=\gamma{\mathcal{H}}_{k,n}+w_{k,k,n+1}(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{Y}_{k,n+1})$ (23)
@@
On a high level, for updating the coefficients in the online GAMLSS, we do the following steps in the outer cycle: we initialise the $\boldsymbol{\eta}_{k,n+1}^{[0,0]}=g_{k}\left(\boldsymbol{\beta}_{k,n}\boldsymbol{X}^{\top}_{k,n+1}\right)$ and this allows us to evaluate $\boldsymbol{u}_{k,n+1}^{[0,0]}$, $\boldsymbol{W}_{k,k,n+1}^{[0,0]}$ and $\boldsymbol{z}_{k,n+1}^{[0,0]}$. From there, we can update the weighted, exponentially discounted Gramian matrices ${\mathcal{G}}_{k,n}\rightarrow{\mathcal{G}}_{k,n+1}^{[0,0]}$ and ${\mathcal{H}}_{k,n}\rightarrow{\mathcal{H}}_{k,n+1}^{[0,0]}$ and subsequently run the online coordinate descent (OCD) algorithm to update the coefficient path in the inner iteration. As in the batch algorithm (see Algorithm 1), we run the outer and inner cycles until convergence.
Input: $\boldsymbol{Y}_{n+1},\boldsymbol{X}_{k,n+1}$ and stored Gramian ${\mathcal{G}}_{k,n},{\mathcal{H}}_{k,n}$, distribution $\mathcal{D}$
1 Initialise the fitted values $\hat{\theta}_{k,n+1}^{[0,0]}=\hat{\boldsymbol{\beta}}_{k,n}\boldsymbol{X}^{\top}_{k,n+1}$ for $k=1,...,p$.
2 Evaluate the linear predictors ${\hat{\eta}^{[0,0]}_{k,n+1}=g_{k}(\hat{\theta}_{k,n+1}^{[0,0]})}$ for $k=0,...,p$.
3 for $m=0,...$ until convergence do
4       forall $k=1,...,p$ do
5             Start the inner cycle.
6             for $r=0,1,...$ do
7                   Evaluate $u_{n+1,k}^{[m,r]}$, $w_{k,k,n+1}^{[m,r]}$ and $z_{n+1,k}^{[m,r]}$ using Equations 4, 5 and 6.
@@
8                   Update ${\mathcal{G}}_{k,n}\rightarrow{\mathcal{G}}_{k,n+1}^{[m,r]}$ and ${\mathcal{H}}_{k,n}\rightarrow{\mathcal{H}}_{k,n+1}^{[m,r]}$ by taking the working vector $z_{n+1,k}^{[m,r]}$ as response variable and weights $w_{k,k,n+1}^{[m,r]}$:
${\mathcal{G}}_{k,n+1}^{[m,r]}$ $=\gamma{\mathcal{G}}_{k,n}+w_{k,k,n+1}^{[m,r]}\left(\boldsymbol{X}^{\top}_{k,n+1}\boldsymbol{X}_{k,n+1}\right)$
${\mathcal{H}}_{k,n+1}^{[m,r]}$ $=\gamma{\mathcal{H}}_{k,n}+w_{k,k,n+1}^{[m,r]}\left(\boldsymbol{X}^{\top}_{k,n+1}z_{n+1,k}^{[m,r]}\right)$
@@
9                  Update $\hat{\boldsymbol{\beta}}_{n,k,\lambda}\rightarrow\hat{\boldsymbol{\beta}}_{n+1,k,\lambda}^{[m,r+1]}$ using OCD based on ${\mathcal{G}}_{k,n+1}^{[m,r]}$ and ${\mathcal{H}}_{k,n+1}^{[m,r]}$
10                   Select the optimal $\lambda$ using IC and set $\hat{\boldsymbol{\beta}}_{n+1,k}^{[m,r+1]}=\hat{\boldsymbol{\beta}}_{n+1,k,\lambda^{\text{opt}}}^{[m,r+1]}$.
11                   Calculate the updated $\hat{\boldsymbol{\eta}}_{k,n+1}^{[m,r+1]}$ and $\hat{\boldsymbol{\theta}}_{k,n+1}^{[m,r+1]}$
12                   Evaluate the convergence.
13
14            End the inner cycle on the convergence of $\hat{\boldsymbol{\beta}}_{k,n+1}^{[m,...]}$
15             Set $\hat{\boldsymbol{\beta}}_{k,n+1}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\beta}}_{k,n+1}^{[m,...]}$, set $\hat{\boldsymbol{\eta}}_{k,n+1}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\eta}}_{k,n+1}^{[m,...]}$ and set $\hat{\boldsymbol{\theta}}_{k,n+1}^{[m+1,0]}\leftarrow\hat{\boldsymbol{\theta}}_{k,n+1}^{[m,...]}$.
16      End the outer cycle if the change in the penalized likelihood is sufficiently small.
17
Output: $\widehat{\boldsymbol{\beta}}_{k,+++n+1}$ and $\widehat{\boldsymbol{\theta}}_{k,n+1}$
Algorithm 4 Online regularized GAMLSS.
Detailed Algorithm
The following paragraph describes the coefficient update for the online GAMLSS algorithm in detail. Algorithm 4 gives an overview. We keep the structure from Section 2.1. To make the discussion clearer, we call an update step a full update for all distribution parameters $\theta_{k,n}\rightarrow\theta_{k,n+1}$ and associated regression coefficients $\boldsymbol{\beta}_{k,n}\rightarrow\boldsymbol{\beta}_{k,n+1}$ for $k=1,...,p$ while after receiving data $\boldsymbol{Y}_{k,n+1}$ and $\boldsymbol{X}_{k,n+1}$, while step might refer to any arbitrary step in the algorithm. A few implementation details in the online GAMLSS are worth discussion:
@@
  • Note that in Line 8, in each inner iteration, we start at the Gramian matrices of the previous full fit, not at the previous iterations $m$ or $r$ of the algorithm, since this would imply adding the $\boldsymbol{X}_{k,n+1}$ and $z_{n+1,k}^{[m,r]}$ multiple times to the Gramian matrices within one update step. However, we can (and should) warm-start the OCD algorithm using the coefficient path from the previous iterations within each inner iteration of the update step.
  • For the online models selection based on the information criteria, we update the exponentially discounted weighted sum of squared residuals to approximate the likelihood (see Equation 21) by taking
    $\text{RSS}_{k,n+1,\lambda}=\dfrac{\left(w_{k,k,n+1}\left(z_{n+1,k}^{[m,r]}-\hat{\boldsymbol{\beta}}_{n,k,\lambda}^{[m,r+1]}\boldsymbol{X}^{\top}_{k,n+1}\right)\right)^{2}+(1-\varphi)\omega_{k,k,n}\text{RSS}_{n,k,\lambda}}{\omega_{k,k,n+1}}$ (24)
    where $\omega_{k,k,n}$ is the mean of discounted weights $\omega_{k,k,n}=1/\tilde{N}\sum_{i=0}^{N}\gamma^{N-i}w_{k,k,i}$ up to $n$, which can easily be tracked online. Subsequently, we select the
    $\lambda^{\text{opt}}_{k,n+1}=\operatorname{arg\;min}_{\lambda}\left(\operatorname{GIC}(\hat{{\mathcal{L}}}_{\lambda},\nu)\right)$
    and update the likelihood based on the RSS by Equation 21.
@@
The main convergence criteria is the difference between the log-likelihood under the distribution $\mathcal{D}$, given the inner and outer iterations current fitted values $\hat{\boldsymbol{\eta}}^{[m,r]}_{k,n+1}$, both for the batch and online case. We track the (exponentially discounted) log-likelihood for checking the convergence in the online case. Additionally, we introduce a breaking criterion if the residual sum of squares increases during the inner cycle by more than a factor of $\varepsilon_{\text{RSS}}$ of the previous inner cycle’s RSS. This criterion can reduce the risk of divergence, while, on the flip-side, ’forces’ the algorithm into a local minima with respect towards one distribution parameter. We observe that divergence is a mainly problem if the chosen parametric distribution assumption does not fit the observed data well, especially in the case of outliers.
@@
  • For the selection of the regularization, we suggest selecting a higher regularization for higher $k$ to avoid overfitting in modelling the conditional scale, kurtosis and skewness (Ziel, 2022). We also note that the $\varepsilon_{\lambda,k}$ for higher moments needs to be smaller, otherwise the lower end of the coefficient path might not approach the OLS solution. Groll et al. (2019) analyse the impact of different shrinkage parameters on batch distributional regression. Also Marcjasz et al. (2023) see the advantages of different regularizations for the distribution parameters when using distributional neural networks for time series forecasting.
Relationship to batch setting
@@
Finally, let us note an important relationship to the batch setting of our proposed algorithm. In the batch setting with repeated fits of the GAMLSS, the weight matrix ${\mathcal{W}}_{k,k}^{[m,r]}=\operatorname{diag}\left(w_{k,k,0},...,w_{k,k,n}\right)$ is updated in every iteration $m,r$ and again, in the next batch fit for $n+1$, the weight matrix is updated for all for $w_{k,k,i}$ from $i=0,...,n+1$ in all iterations. In the online setting, ${\mathcal{W}}_{k,k,n}$ is the weight matrix after convergence of update step $n-1\rightarrow n$. In the update step $n\rightarrow n+1$, we cannot update ${\mathcal{W}}_{k,k,n}$ anymore (see also the Assumptions noted in Section 2.2). Therefore, we set
${\mathcal{W}}_{k,k,n+1}^{[m,r]}=\begin{pmatrix}(1-\varphi){\mathcal{W}}_{k,k,n}&0\\ 0&w_{k,k,n+1}^{[m,r]}\end{pmatrix}$ (25)